Fermi National Laboratory
 

Accelerator Update for January 10, 2007 - January 12, 2007

 

Wednesday January 10

The day shift began with Operations monitoring store 5172 (183.36E30) and a Recycler stash (12.9E10), with the Antiproton Source (Pbar) stacking (77.4E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off due to absorber work, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

A Pbar expert reported at 8:19 AM that the extraction shutter had been repaired.

At 8:50 AM, the Tevatron Electron Lens (TEL1) tripped off due to a vacuum problem. It reset okay.

At 8:52 AM, a Linac Klystron RF station (KRF6) tripped off due to a LCW flow problem. Operators reset the station.

At 9:30 AM, a Linac RF station (LRF3) had its breaker trip. Linac technicians reset the breaker.

At 10:49 AM, five MI RF stations tripped off on a power amplifier power supply fault. Of the five stations, MIRF14 wouldn’t reset. An expert investigated and reset the station.

At 11:19 AM, the MI South Anode power supply tripped off on a current monitor indication. It reset.

A Pbar expert began a stacktail study at 11:38 AM; this study will disrupt the beam.

At 11:41 AM, Operations disabled beam so that Preaccelerator experts could access the H- Source. Power supply experts brought a MI power supply back on line during this time.

The MI power supply work was completed at 11:57 AM.

The H- Source work was completed at 12:07 PM.

The Linac work was completed at 12:32 PM. NuMI resumed taking beam.

A MI expert began a slipstacking study at 1:07 PM.

At 2:59 PM, a Booster RF station (BRF6) tripped off due to sparking. It reset.

At 3:30 PM, KRF3 tripped off for the first of two times in thirteen minutes. The trip was due to a LCW flow regulation problem. Operations had some trouble getting the station back up after the second trip. Beam resumed at 4:03 PM.

At 4:22 PM, Operations reported that KRF3 was not phase regulating. Linac experts reported that the mixing valve for the cavity water skid was the problem. At 5:22 PM, the experts reported that a broken flow turbine on the water skid was the real problem. An expert was called in to make repairs.

At 6:29 PM, BRF13 & 14 tripped off due to a vacuum pump trip. Everything reset okay.

Linac experts reported at 7:47 PM that KRF3 had been repaired. NuMI resumed taking beam. Stacking didn’t resume right away; it took several resets. Stacking resumed at 7:59 PM.

At 8:35 PM, BRF6 tripped off for the first of two times in thirty minutes due to sparking.

At 9:24 PM, the NuMI beam permit tripped off due to a problem with beam positions. This was due to a H- Source glitch. The permit reset.

The Recycler completed its last of two stashes around midnight.

Thursday January 11

The midnight shift began with Operations monitoring store 5171 (51.46E10) and the stash (195.6E10), with Pbar stacking (10.7E10), and with NuMI taking beam.

At 1:27 AM, the Pbar beam permit tripped off due to a current drop on a power supply (HV102).

The Recycler completed its last of three stashes for the midnight shift at 5:19 AM.

At 5:30 AM, D0 reported that a crate power supply for their calorimeter had tripped off. They will need a 3-4 hour access to repair a failed fan and the power supply.

A TeV expert began an end-of-store study at 6:26 AM.

Operations terminated store 5171 at 7:06 AM.

Operations had CDF and D0 prepared for access by 7:20 AM.

An interlock expert reported at 8:03 AM that the Meson safety system tests were complete.

At 8:51 AM, KRF5 tripped off due to a spark. It reset.

At 10:07 AM, a TeV RF station (TRF8) tripped off. An expert investigated and found nothing wrong. He turned the station on.

Operations put the TeV through a dry squeeze at 11:20 AM.

Operations put the TeV into shot setup at 11:51 AM.

Operations established store 5174 at 1:46 PM with an initial luminosity of 245.9E30.

The Recycler completed the last of two transfers during the day shift at 2:08 PM.

A MI expert began a 4.5-hour slipstacking study at 2:21 PM.

A MI expert began a MI-8 collimator study at 2:49 PM.

At 3:53 PM, the MI South anode power supply tripped off for the first of two times in about 30 minutes on an AC overload function. Operators reset it after the first trip, but contacted an expert after the second trip. The expert reported at 4:31 PM that the trips were due to a faulty current monitor in the yard. Operations will lower the intensity until the monitor can get repaired, which would be a 12-hour repair. A tweak to the monitor could stop the trips for a while. This work would take about 30 minutes.

Operations reported at 4:41 PM that operators had searched and secured the Meson enclosures M02, 3, 4, and 5.

At 10:58 PM, TRF8 tripped off again.

The Recycler completed the last of three stashes at 11:10 PM.

Friday January 12

The midnight shift began with Operations monitoring store 5174 (73.67E30) and the stash (200.4E10), with Pbar stacking (29.9E10), and with NuMI taking beam.

At 1:52 AM, Operations reported that a Booster vacuum pump (B:IPQ16) had tripped off three times in the last 20 minutes. Operators left the pump off for a couple of hours and then turned it back on.

At 1:57 AM, the P2 transfer line beam permit tripped off due to the current drift of a power supply. NuMI and MI also tripped. Everything reset.

At 2:43 AM, the Pbar beam permit tripped off due to a power supply trip (HV100). Operators investigated and found a bad SCR indication. The supply tripped again at 2:54 AM. Operations reset the supply.

At 3:14 AM, LRF2 tripped off. It reset okay.

By 5:25 AM, HV100 had tripped off two more times. Operations called in two power supply experts and put Pbar in standby. The experts reported at 6:24 AM that they had found and replaced two bad SCR firing modules. Stacking resumed.

At 6:51 AM, a Debuncher RF station (DRF-1) was missing its output voltage. An expert investigated.

At 7:57 AM, a water expert reported that the TeV had lost 300 gallons of LCW. Water technicians will search the TeV service buildings and CUB for the leak.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will consider allowing thirty minutes for the MI current monitor tweak. Meson experts will take beam to the M01 absorber for a few hours on Saturday and Sunday.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/12/2007    email Fermilab